The temperature, increased by 65 degree, is the same as 5 times the temperature
decreased by 15 degree. Find the temperature

Answer by nomad_55(6)

t+65=5t-15
+15 +15
t+80=5t
-t -t

80=4t
/4 /4

20=t

(20)+65=5(20)-15

85=100-15

85=85
good!